Week of 171204

WLCG Operations Call details

  • At CERN the meeting room is 513 R-068.

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Web
  • Whenever a particular topic needs to be discussed at the daily meeting requiring information from site or experiments, it is highly recommended to announce it by email to wlcg-operations@cernSPAMNOTNOSPAMPLEASE.ch to make sure that the relevant parties have the time to collect the required information or invite the right people at the meeting.

Tier-1 downtimes

Experiments may experience problems if two or more of their Tier-1 sites are inaccessible at the same time. Therefore Tier-1 sites should do their best to avoid scheduling a downtime classified as "outage" in a time slot overlapping with an "outage" downtime already declared by another Tier-1 site supporting the same VO(s). The following procedure is recommended:

  1. A Tier-1 should check the downtimes calendar to see if another Tier-1 has already an "outage" downtime in the desired time slot.
  2. If there is a conflict, another time slot should be chosen.
  3. In case stronger constraints cannot allow to choose another time slot, the Tier-1 will point out the existence of the conflict to the SCOD mailing list and at the next WLCG operations call, to discuss it with the representatives of the experiments involved and the other Tier-1.

As an additional precaution, the SCOD will check the downtimes calendar for Tier-1 "outage" downtime conflicts at least once during his/her shift, for the current and the following two weeks; in case a conflict is found, it will be discussed at the next operations call, or offline if at least one relevant experiment or site contact is absent.

Links to Tier-1 downtimes

ALICE ATLAS CMS LHCB
  BNL FNAL  

Monday

Attendance:

  • local: Julia (WLCG), Maarten (ALICE, WLCG), Ivan (ATLAS), Gavin (Comp), Alberto (monitoring), Andrea M (MW), Andrei (DB), Marian (network), Roberto (storage), Yolanda (storage), Kate (DB, chair)
  • remote: Antonio F (LHCb, CNAF), Andrew P (NL-T1), Antonio Y (PIC), Stephan (CMS), Xavier (KIT), Kyle (OSG), David (IN2P3), Dave M (FNAL), Di Qing (TRIUMF), John K (RAL), Xin (BNL), Jens (NDGF), Chi-Hsun (ASGC)

Experiments round table:

  • ATLAS reports ( raw view) -
    • Stable operation at 350k cores level.
    • HLT off - 50k cores for GRID since 27.11
    • T0 off - 19k cores for GRID since 30.11.2017
    • Slice tests for Christmas data reprocessing - ongoing
    • CNAF data replication 620 TB remaining
    • Grafana job monitoring / HDFS problems - solved.

  • CMS reports ( raw view) -
    • Tier-0 transfer backlog dropped below 5 PB now that data taking stopped, tuning of system continues
    • Tier-0 CPU resources now used by production
    • CMS compute systems busy

  • ALICE -
    • Very high activity on average
      • New records during the weekend: 162k concurrent jobs overall, 84k at CERN
    • CERN: EOS-ALICE not reachable from outside Tue afternoon (GGUS:132120)
    • CERN: CASTOR reconfigured for reco at high scale (50k jobs) - thanks!

  • LHCb reports ( raw view) -
    • Activity
      • Stripping validation, user analysis, MC
    • Site Issues
      • PIC and IN2P3 are not accepting some kind of certificates. Most likely a dCache related issues (GGUS:132237)
        • see the MW Officer report below

Sites / Services round table:

  • ASGC: There will be a downtime from 6/12/2017 2pm UTC to 7/12/2017 2pm UTC
  • BNL: NTR
  • CNAF: Next week we will start working on the tape library recovery and also the installation of the new storage.
  • EGI: nc
  • FNAL: NTR
  • IN2P3: During the maintenance on last Tuesday, we configure one CE to allow ALICE and LHCb jobs to run on CentOS7 WNs. No problem has been observed. ALICE and LHCb are running on SL6 and C7.
  • JINR: NTR
  • KISTI: nc
  • KIT:
    • Rolling updates for dCache as announced last week:
      • Nov 28th, LHCb: Failed to roll out the update gracefully. SE was down from 9:30 to 11:30 a.m.
      • Nov 29th, CMS: Update went well w/o major problems for clients afaik. Though since the update srm-get-metadata broke and we suggested to switch over to gfal-xattr for querying SRM about a replica's locality.
      • Nov 30th, ATLAS: Happened on Nov 29th by accident. Complete restart of all dCache services meant that the SE wasn't writable for about 5 hours (10:15 till ~15:00 o'clock)
    • LHCb received an increase of 1 PB for their online disk space reservation "LHCb-Disk" (milestone 2017 resources).
  • NDGF: NTR
  • NL-T1: During last Tuesday's maintenance, the SARA dCache was switched back to the dCache internal Zookeeper, as a workaround for the communication issues. We hope this will be stable until the root cause is found.
  • NRC-KI: nc
  • OSG: We were unaware of the GGUS certificate update last Wed. Information flow to be followed up.
  • PIC:
    • New disk pools to cover full 2017 pledge finally arrived at PIC, to be installed and put into service in the coming months
    • ATLAS and LHCb transfers: dcache movers reoptimized and some pools rebooted last week, also gridftp doors and door srmlhcb restarted today
    • CMS request for singularity by Feb 2018, in discussion with VO to clarify if request implies allowing privileged execution, and potential interferences with our SL6 to C7 and PBS to HTCondor migration plans
  • RAL: There are problems with the latest trust anchor release in the UK. We are looking into this.
  • TRIUMF:
    • Tape library robot stuck on Nov 27, however it had been successfully fixed by IBM only on Nov 29.
    • Site wide power cut on Dec 2. The power cut started around 9:40am local time, and got back around 11:50am.

  • CERN computing services: NTR
Privileges on execution of containers to be clarified during Container WG meeting and discussed internally in CMS.
  • CERN storage services: NTR
  • CERN databases
    • ALICE online patching this afternoon
    • LHCB online and PDBR patching tomorrow afternoon
    • CMS (all databases) patching next Monday
    • ATLAS (all databases) patching this week
    • All interventions are expected to be transparent
    • CMSONR ADG issues understood - bug triggered by certain queries - following up with Oracle. CMSSW queries triggering the bug were separated in a separate HW setup.
  • GGUS: NTR
  • Monitoring:

  • MW Officer:
    • issue with UK eScience CA certs since lcg-CA v1.88.1 GGUS:132205. under investigation
      • New certificate signed with SHA2 instead of SHA1; replacement was not as transparent as expected; Java-based services using the Bouncy Castle security module are affected. Some workarounds in place, no conclusion yet on generic solution.
    • Several DPM SRM crashes reported last week by ATLAS sites. They all happened after being contacted by a specific user DN and IP. We identified the issue and a new DPM release is under preparation.
  • Networks: NTR
  • Security: NTR

AOB:

Edit | Attach | Watch | Print version | History: r28 < r27 < r26 < r25 < r24 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r28 - 2017-12-07 - MaartenLitmaath
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback